Gabrielle, a chemistry student has a 70% alcohol solution and 40% alcohol solution. Her teacher instructed her to prepare 150ml of a solution that is 60% alcohol by mixing amounts from each of the available solutions. How much of each solution should she use?
Answer by nerdybill(7384)

Gabrielle, a chemistry student has a 70% alcohol solution and 40% alcohol solution. Her teacher instructed her to prepare 150ml of a solution that is 60% alcohol by mixing amounts from each of the available solutions. How much of each solution should she use?
.
Let x = amount of 70% alcohol solution
then
150-x = amount of 40% alcohol solution
.
.70x + .40(150-x) = .60(150)
.70x + 60 - .40x = .60(150)
.30x + 60 = .60(150)
.30x + 60 = 90
.30x = 30
x = 100 ml of 70% alcohol solution
.
40% alcohol solution:
150-x = 150-100 = 50 ml